### v2

# Anomaly Detection in Complex Dynamical Systems: A Systematic Framework Using Embedding Theory and Physics-Inspired Consistency

Abstract
Anomaly detection in complex dynamical systems is essential for ensuring reliability, safety, and efficiency in industrial and cyber-physical infrastructures. Predictive maintenance helps prevent costly failures, while cybersecurity monitoring has become critical as digitized systems face growing threats. Many of these systems exhibit oscillatory behaviors and bounded motion, requiring anomaly detection methods that capture structured temporal dependencies while adhering to physical consistency principles. In this work, we propose a system-theoretic approach to anomaly detection, grounded in classical embedding theory and physics-inspired consistency principles. We build upon the Fractal Whitney Embedding Prevalence Theorem that extends traditional embedding techniques to complex system dynamics. Additionally, we introduce state-derivative pairs as an embedding strategy to capture system evolution. To enforce temporal coherence, we develop a Temporal Differential Consistency Autoencoder (TDC-AE), incorporating a TDC-Loss that aligns the approximated derivatives of latent variables with their dynamic representations. We evaluate our method on the C-MAPSS dataset, a benchmark for turbofan aeroengine degradation. TDC-AE outperforms LSTMs and Transformers while achieving a nearly 200x reduction in MAC operations, making it particularly suited for lightweight edge computing. Our findings support the hypothesis that anomalies disrupt stable system dynamics, providing a robust signal for anomaly detection.
Index Terms:
Complex Dynamical Systems, Anomaly Detection, System Theory, Embedology, Physics-Informed Machine Learning, Predictive Maintenance, Edge Computing
I Introduction
Anomaly detection in complex physical dynamical systems is a critical research area as industrial and engineered systems become more sophisticated. Identifying deviations from expected behavior is essential for ensuring reliability, safety, and efficiency. This is particularly relevant in predictive maintenance and cybersecurity monitoring. In industrial systems and rotating machinery, early fault detection helps prevent costly failures and downtime [1]. Meanwhile, as critical infrastructures, such as power grids and water distribution systems, become increasingly digitized, the risk of cyber threats has grown significantly [2, 3]. Recent attacks [4, 5] on cyber-physical systems highlight the need for robust monitoring techniques to detect both malicious intrusions and system failures. Ensuring the security and stability of these dynamical systems requires adaptive anomaly detection methods capable of addressing evolving threats and operational challenges.
Many complex dynamical systems exhibit oscillatory behaviors and bounded motion, fundamental characteristics of both natural and engineered processes. The study of such systems intersects with two key fields: time-series modeling and the incorporation of physical laws. Given that many physical systems display structured temporal dependencies, effective modeling requires methods that capture correlations across time. Simultaneously, classical physics, which governs tangible objects and engineered systems, is largely defined by causal and deterministic principles, often described through differential equations and conservation laws. This study aims to bridge time-series modeling with physics-inspired approaches to develop more effective and sustainable anomaly detection methods.
II Related Work
II-A Time-Series Modeling
Models like LSTMs and Transformers with attention mechanisms have been successfully applied to time-series anomaly detection in fields such as mechanical engineered system, aerospace, and industrial monitoring [6, 7, 8]. While these methods effectively capture long-range dependencies and irregular temporal patterns, they are computationally expensive. For example, evaluated on common benchmark datasets, sequence lengths up to 500 time steps were required to achieve competitive performance [9]. The high memory requirements often exceed the constraints of typical MCU-level devices, making real-time deployment challenging. For example, a typical LSTM with 500 time steps and 50 hidden units already surpasses the available memory of common low-power devices, limiting its practical applicability. Additionally, the sequential nature of RNNs restricts parallelization, further increasing computational cost [10]. These challenges highlight the need for alternative approaches that balance computational efficiency with robust anomaly detection, aligning with broader goals of sustainability and practical deployability.
II-B Physics-Informed Methods
In many applications, incorporating domain knowledge can help reduce computational demands. Physics-informed neural networks (PINNs) embed physical laws directly into neural networks, enabling them to leverage known system dynamics [11, 12, 13]. However, applying PINNs to complex dynamical physical systems remains challenging. These systems involve numerous interdependent physical principles, and explicitly modeling them within a neural network would require domain expertise, making large-scale applications impractical [14, 15].
II-C Benchmark Use Case: Turbofan Aeroengine Degradation
We use the C-MAPSS dataset as a benchmark for studying complex dynamical systems due to its realistic representation of turbofan aeroengine degradation. Aeroengines are complex dynamical systems governed by nonlinear, time-dependent interactions of physical processes, making them an ideal test case for evaluating anomaly detection methods in real-world settings.
In the literature, anomaly detection for the C-MAPSS dataset generally follows two main approaches. The first focuses on fleet-level anomaly detection, where engines with shorter lifetimes are classified as abnormal based on their total life cycles [16, 17]. This method aims to distinguish early failures from normal operating conditions at a system-wide level.
The second approach considers individual engine degradation, defining anomalies based on a 60/40 time-based split [18, 19]. In this setup, the first 60% of an engine’s life is labeled as normal, while the final 40% is considered abnormal. The dataset is then divided into train/test subsets, and models are evaluated based on their ability to classify each time step accordingly.
Since we aim to develop methods that capture system dynamics, we consider the second approach a more suitable benchmark, as it focuses on time-dependent degradation rather than static fleet-level classification.
Previous work on anomaly detection in this setting has explored various deep learning models. One study employed an LSTM-based approach [19], leveraging recurrent structures to model time dependencies. Another approach used a standard autoencoder (AE) without explicit temporal modeling [18]. More recent research has investigated Transformer-based models, which excel at capturing long-range dependencies but introduce high computational costs and require extensive training data due to their large parameter space [20, 7]. Moreover, both Transformer studies formulated the problem as a multiclass classification task, with one using a slightly different dataset and the other applying the method directly to C-MAPSS. However, defining well-separated fault categories in a complex dynamical system is challenging in practice, as non-trivial interactions between multiple physical components create highly unpredictable behaviors. As system dynamics grow more intricate, these interactions become even less predictable, further complicating precise fault categorization.
II-D Proposed Approach
Instead of relying on complex models to differentiate normal and abnormal behavior, we adopt a system-theoretic approach inspired by classical embedding theory to analyze the system’s latent dimensionality. By incorporating physics-inspired consistency principles, we approximate the underlying causal mechanisms without explicitly enforcing physical laws. We hypothesize that in a stable regime, complex systems exhibit predictable behavior, enabling well-approximated low-dimensional embeddings. In contrast, anomalies introduce additional complexity, disrupting these stable relationships. We aim to leverage this property to detect abnormal states when the learned embedding no longer adequately captures the system dynamics.
III Theoretical Foundation & Methods
We study the mathematical foundations of embedding bounded dynamical systems into an autoencoder’s latent space, aiming to ensure reliable mapping of system dynamics for anomaly detection and health monitoring.
Consider for example an oscillatory dynamical system, specifically a damped-driven pendulum. Initially, the system remains within a bounded oscillatory regime; however, due to the influence of drift terms, it gradually deviates, eventually departing from the confined state space. This perturbation effect is illustrated in Figure 1, where the phase space trajectory initially remains within a bounded region but progressively transitions into an unbounded state. This formulation highlights that the state variables of an oscillatory system, namely the angular displacement $\theta$ and its derivative $\dot{\theta}$, define a phase space where the system’s behavior can be systematically analyzed. Within this phase space, bounded regions correspond to normal oscillations, where the system follows a predictable, periodic trajectory. However, once the system leaves this bounded region, it signifies the onset of unbounded oscillations or deviations from regular behavior. For formal definitions and details on the system’s implementation, refer to Appendix A.
Figure 1: Phase plot of a damped forced motion, showing $\theta$ and $\dot{\theta}$. The blue trajectory represents the stable phase space evolution, while the red trajectory highlights an unstable deviation from the stable state.
We explore whether this concept extends to complex dynamical systems and the constraints and conditions for a successful transition. The challenge lies in preserving phase space properties in high-dimensional, nonlinear systems while maintaining topological consistency and dynamical coherence. To address this, we examine the mathematical and computational requirements for embedding system dynamics into a reduced representation, guiding the design of effective autoencoder-based anomaly detection.
III-A Embedding Complex System Dynamics into a Latent Space
We begin by formalizing the mathematical foundation for embedding high-dimensional dynamical systems into a reduced representation. Consider a complex bounded dynamical system with state variables $x\in\mathbb{R}^{m}$, where $m$ represents the dimension of the physical system and is typically much larger than the number of available measurements $(m\gg k$).
At the measurement level, we access observables through a measurement function:
$\mu:\mathbb{R}^{m}\to\mathbb{R}^{k},$ (1)
where $\mu$ extracts a lower-dimensional set of measurements $y=\mu(x)$, with $y\in\mathbb{R}^{k}$, forming the input to the autoencoder.
The encoding level is defined by:
$\mathcal{E}:\mathbb{R}^{k}\to\mathbb{R}^{n},$ (2)
where $\mathcal{E}$ maps the measurement space to an embedding space of dimension $n$.
We analyze the latent space encoding of a dynamical system from a systems-theoretic perspective. Ideally, such systems in a stable state evolve on smooth manifolds, but real-world dynamics often unfold on complex, non-smooth sets due to measurement noise and the inherent need for machines to exhibit more intricate dynamics to perform their functions effectively. A key aspect of this generalization is the notion of dimensionality for compact subsets, which leads to the definition of box-counting dimension.
Definition 1 (Box-Counting Dimension [21]).
Let $A$ be a bounded subset of $\mathbb{R}^{k}$. For each $\epsilon>0$, let $N_{\epsilon}(A)$ denote the smallest number of sets of diameter at most $\epsilon$ required to cover $A$.
The box-counting dimension of $A$ is defined as
$\dim_{B}(A)=\lim_{\epsilon\to 0}\frac{\ln N_{\epsilon}(A)}{\ln(1/\epsilon)}$
provided the limit exists.
In this context, $\epsilon$ provides a measure of the complexity exhibited by motion beyond smooth manifold assumptions.
Equipped with this notion of dimensionality, we now state the embedding theorem that generalizes classical results to fractal structures.
Theorem 1 (Fractal Whitney Embedding Prevalence Theorem [22]).
Let $A$ be a compact subset of $\mathbb{R}^{k}$ with box-counting dimension $d$, and let $n$ be an integer greater than $2d$. For almost every smooth map $F:\mathbb{R}^{k}\to\mathbb{R}^{n}$,
  1. 1.
    $F$ is one-to-one on $A$.
  2. 2.
    $F$ is an immersion on each compact subset $C$ of a smooth manifold contained in $A$.
We discuss the necessary assumptions to apply this formalism to our system representation.
Compactness of the System State: We assume that the key drivers of the physical process form a bounded and closed subset of $\mathbb{R}^{k}$, ensuring compactness. This assumption reflects the practical reality that real-world physical systems with inherently bounded dynamics, when operating in a stable state, remain confined within a finite region of their state space.
Informative Measurement Function: We assume that the measurement function retains sufficient information about the system dynamics. Specifically, it must be injective enough over $A$ to avoid collapsing distinct states, ensuring that the measurement space captures the necessary complexity of the system.
Dimensionality of the System: We assume that the measurement-level set $A\subset\mathbb{R}^{m}$ has a well-defined box-counting dimension $d$, which reflects the effective complexity of the system’s dynamics. This dimensionality determines the necessary embedding space.
Under these conditions, the theorem is applicable. This allows us to estimate the required embedding dimension $n$ to achieve a faithful description of the system’s dynamics.
After analyzing the necessary embedding dimension, we now turn to the type of embedding that best captures the dynamics of the system. We draw inspiration from simple dynamical systems and propose using state-derivative pairs as an embedding strategy. This choice is motivated by two key principles. Physical laws typically encode causal relationships between a system’s state and its derivatives, making this a natural way to describe a dynamical process. When chosen properly, state-derivative pairs help reduce redundancy by maintaining linear independence between embedding dimensions.
To effectively utilize this type of embedding, we need more than just a one-to-one mapping—we also require consistent differential structure. In other words, small changes in the latent space should produce meaningful, non-degenerate changes in the measurement space. A full-rank Jacobian is a necessary and sufficient condition to guarantee this kind of well-behaved, locally reversible mapping [23].
We hypothesize that leveraging this type of embedding allows for two key insights:
  • Deviations beyond a certain region associated with normal variability indicate anomalous behavior, as suggested by phase space considerations.
  • As the dynamics become more complex, this leads to an increase in the box-counting dimension. By Th. 1, this increase suggests that the chosen embedding dimension is insufficient to fully describe the system. Consequently, we hypothesize that this insufficiency results in a collapse of the dynamical description, leading to the loss of the causal approximation between the state and its derivative.
III-B Application to Autoencoder-based Representations
With the theoretical framework established, we apply it to a practical case in two steps. First, we select a stable, bounded dynamical system with accessible measurements. For this, we use the NASA C-MAPSS dataset, which provides time-series sensor data from aircraft engines. Second, we develop an autoencoder that learns embeddings aligned with our framework, ensuring a one-to-one mapping and immersion within the normal operating regime of the system.
III-B1 C-MAPSS Dataset
The NASA Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset [24] is a widely used benchmark in prognostics and health management studies [25, 26, 27].
The dataset consists of multivariate time series capturing engine operation under varying initial wear, manufacturing differences, and noisy sensor readings. Each engine starts in a normal state before gradually developing a fault.
While originally designed for Remaining Useful Life (RUL) estimation, the C-MAPSS dataset has become more popular lately in health management studies for anomaly detection, fault diagnostics, and predictive maintenance [16]. Its structured time-series format and realistic degradation patterns make it valuable for developing and testing machine learning models aimed at early fault detection and system health assessment.
We use the FD001 subset of the C-MAPSS dataset, which contains run-to-failure time series from 100 independent engines operating under a single operating condition but exhibiting varying fault progressions. This substantial number of individual systems allows us to assess the generalization capability of our method across multiple degradation trajectories. For anomaly detection, we follow a 60/40 split, where the first 60% of cycles are labeled as normal and the remaining 40% as anomalous, as suggested in similar studies [18, 19]. A random 80/20 train-test split based on separate engines ensures robust evaluation.
The dataset includes sensor readings for temperature, pressure, fan and core speed, efficiency ratios, fuel flow, and coolant bleed parameters, providing insights into engine performance and degradation.
III-B2 Applying the Theoretical Foundation
The NASA C-MAPSS dataset provides a measurement space of dimension $k=24$, representing sensor readings from the turbine system. To determine a suitable embedding dimension in accordance with Theorem 1, we establish the following assumptions about the system’s essential dynamics.
As argued in the theory section, the turbine operates in a stable regime, where its state evolution remains within a well-defined, physically constrained region. This ensures that the system’s key drivers form a bounded and closed subset of $\mathbb{R}^{k}$, thereby satisfying the compactness assumption.
We hypothesize that the turbine’s core behavior is governed by three fundamental state-derivative pairs, corresponding to rotation, temperature, and pressure along with their respective rates of change. These represent the system’s intrinsic dynamics and serve as a basis for estimating its fractal dimension. In Appendix B, we computed the box-counting dimension numerically for the attractor of a damped driven pendulum, yielding $d\approx 1.3$. For a perfect mathematical pendulum, this value reduces to $d=1$, reflecting its simple periodic structure (a circle). Given the structural similarities between the fundamental dynamics of the pendulum and the turbine’s primary state variables, we assume a similar level of dynamic complexity here. However, this framework provides room to incorporate domain knowledge, such as machine tolerances and operational constraints, to refine this assumption further. This leads to an estimated box-counting dimension $d<4$ for the combined system. Applying Theorem 1, this result suggests the selection of an embedding dimension of $n=8$.
The measurement function in the C-MAPSS dataset is assumed to retain sufficient information about the system’s essential dynamics. Given the structured nature of the sensor readings, it provides a reliable representation of the turbine’s state evolution without collapsing distinct states, thereby satisfying the injectivity condition.
Based on our approach, we select four state-derivative pairs for the embedding. To ensure that this representation is at an immersion in the subspace corresponding to the system’s normal or stable regime, we apply a numerical method to verify whether the Jacobian has full rank in the normal state of the test data. While this does not constitute a strict mathematical proof, it serves as a reliable guideline for assessing the validity of the embedding. SVD is particularly powerful for analyzing the rank of a matrix. The rank of $X$ can be determined by counting the number of non-zero singular values in $\Sigma$. If all singular values are non-zero, the matrix is full rank [28].
Theorem 1 assumes a smooth map and that it applies to ”almost every” map. A $\tanh$-activated neural network encoder reasonably satisfies the smoothness assumption. The ”almost every” condition applies to all smooth maps, whereas neural networks form a restricted subset. However, by universal approximation, a sufficiently large network can approximate such an embedding, making the theorem a strong and relevant practical guideline.
III-B3 Temporal Differential Consistency informed Autoencoder
As a foundational step, we require a method to approximate the first-time derivative. A widely used approach is the central difference method, which estimates the derivative of a function by computing the slope between two points symmetrically positioned around the point of interest. The central difference formula for the first derivative of a function $f(t)$ is:
$\dot{f}(t)\approx\frac{f(t+\Delta t)-f(t-\Delta t)}{2\Delta t}$
where $\Delta t$ represents a small step size. This method is often preferred over forward or backward differences due to its higher accuracy ($O(\Delta t^{2})$), meaning that the approximation error decreases quadratically as $\Delta t$ decreases [29].
To accurately capture temporal dynamics, we incorporate the central difference method into the training framework of the latent space. The latent representations at the previous ($t-1$) and next ($t+1$) time steps are used to approximate the first-time derivative of the static latent variables ($z$) using the central difference formula. This derivative is then used as a target for the dynamic latent variables ($\dot{z}$), ensuring consistency with the central difference approximation, scaled by the time interval $\Delta t$. A schematic representation of this approach is shown in Fig. 7 in Appendix C.
To incorporate temporal dynamics into the training process, we introduce a temporal differential consistency loss (TDC-Loss). This loss enforces consistency between the approximated time derivative of the static latent variables ($z$) and the output of the corresponding dynamic latent variables ($\dot{z}$). During training, the central difference method is used to estimate the time derivative of $z$, which is then compared with $\dot{z}$. The TDC loss serves as a regularizer for the standard reconstruction loss, ensuring that the autoencoder learns a latent representation that captures both state and derivative information. A compact version of the pseudocode for TDC-informed training is presented in Algo. 1 in Appendix B. We use a highly compact autoencoder with three hidden layers in both the encoder and decoder. For details on implementation and hardware, see Appendix B.
III-B4 Consistency Metrics
We employ an embedding based on state-derivative pairs, where the derivative is approximated using the central difference method. While numerical differentiation is generally ill-posed, particularly in the presence of noise [30], our approach does not seek a fully precise description of the system’s dynamics. Instead, we aim for a causal approximation that reliably indicates when the system deviates from normal variability. By choosing a sufficiently large embedding dimension, we ensure that the system operates on a simple attractor geometry. This allows us to expect a nearly constant trendline for both the state and its derivative. To evaluate this, we introduce two metrics
To ensure a balanced causal approximation between state and derivative, their variations must remain comparable. We define this using the ratio of their min-max scaled variances:
$\eta=\frac{\sigma_{\dot{x}_{\text{scaled}}}^{2}}{\sigma_{x_{\text{scaled}}}^{2}},$ (3)
where $\sigma^{2}$ denotes variance after min-max scaling. If $\eta\approx 1$, the variations are well-matched, making central difference approximation suitable. When $\eta\ll 1$, the derivative is overly smoothed, potentially suppressing meaningful dynamics. Conversely, $\eta\gg 1$ indicates excessive noise in the derivative, requiring additional smoothing or refined differentiation techniques.
The second metric evaluates how well the approximated derivative maintains consistency with the state transitions. Specifically, we compute the mean squared error (MSE) between the integrated derivative and the actual state difference over small time intervals:
$\rho=\frac{1}{N}\sum_{i=1}^{N}\left(x_{i}-\sum_{j=1}^{i}\dot{x}_{j}\Delta t\right)^{2}.$ (4)
This metric does not require knowledge of the true derivative but serves as a self-consistency check. By integrating the derivative, short-term fluctuations are smoothed, making long-term trends more apparent. A significant increase in this error in anomalous conditions suggests a breakdown in the causal approximation, indicating that the embedding dimension is insufficient to describe the system dynamics adequately.
III-C Anomaly Detection Logic
We integrate the developed mathematical foundations into anomaly detection by leveraging state-derivative pairs in the latent space. Normal states are assumed to lie within a confined region, with deviations indicating anomalies.
Training uses the first 60% of each engine’s data, with an 80/20 split into training and validation sets. Thresholds for each latent node are set as percentiles of the validation data and optimized using a classification metric (e.g., F1-score). This optimization can be performed based on the classification performance on the validation dataset. The threshold remains fixed across all engines. During testing, the model infers the derivative representation from a single time step, as measurements at time $t$ inherently encode rate changes.
Detection logic:
  • If all latent nodes exceed their thresholds during normal operation, it is classified as a false positive (FP); otherwise, it is a true negative (TN).
  • If at least one latent node exceeds its threshold in the anomaly range, it is classified as a true positive (TP); otherwise, it is a false negative (FN).
Detection rules can be adapted based on specific use cases and risk strategies, such as requiring multiple nodes to exceed thresholds before flagging an anomaly.
IV Results
IV-A Consistency of the Proposed Approach
First, we examine the consistency of the proposed approach. The loss terms consistently converge over 50 training epochs across five independent runs, indicating stable training behavior. Both the standard Mean Squared Error (MSE) loss and the newly introduced TDC loss term exhibit this convergence. For a visual representation, refer to Fig. 6 in Appendix C. We conducted a Jacobian rank analysis using Singular Value Decomposition (SVD). Applying a zero-threshold of $\epsilon=10^{-9}$, all samples in the test dataset exhibit a full-rank Jacobian. This result suggests that the mapping $\mathcal{E}$, which projects the measurement space $\mathbb{R}^{k}$ to an embedding space $\mathbb{R}^{n}$, behaves as an immersion, preserving the differential structure.
The first metric, $\sigma$, confirmed the validity of the central difference method for approximating the first time derivative, with values ranging from 0.7 to 1.2 across engines in the training split. The second metric assessed the correlation of state derivatives, capturing causal relationships under normal and anomalous conditions. Table I summarizes the mean $\pm$ standard deviation of the correlation coefficient $\rho$ in both cases.
TABLE I: Comparison of mean $\pm$ standard deviation values of $\rho$ for normal and anomalous conditions across latent node pairs.
Latent Node Pair Normal Condition ($\rho$) Anomalous Condition ($\rho$)
(0-4) $0.00037\pm 0.00012$ $0.00044\pm 0.00016$
(1-5) $0.00019\pm 0.00002$ $0.00034\pm 0.00015$
(2-6) $0.00026\pm 0.00007$ $0.00079\pm 0.00027$
(3-7) $0.00048\pm 0.00014$ $0.00073\pm 0.00011$
The results indicate that in the normal regime, the correlation values remain low with minimal variation. However, in the anomalous regime, we observe a slight increase in mean values and a broader standard deviation range, suggesting a stronger deviation in state derivative relationships under anomaly conditions.
IV-B Phase Space Representation and State-Derivative Relationships
Figure 2: (a) Phase-space plot of the latent node pair (2,6) average. (b) State-derivative consistency metric for this pair. (c) A detailed view of (a) (highlighted grey box), showing phase-space trajectories and the attractor region at the beginning of the system dynamics. The timestep color map applies to all figures.
Fig. 3a-f presents two state-derivative pairs along with their respective phase space representations. A gradual increase or decrease in both state and derivative nodes is observed as the system transitions toward the anomalous range. In the phase space plots, we applied a sliding window of three and slightly smoothed the node values for visualization purposes. This processing was not used in the actual anomaly detection model. During the initial phase of the system dynamics, the states evolve more gradually, supporting the hypothesis that, in the normal regime, the dynamics follow a simple and stable attractor geometry. The phase space representation further indicates that normal states are embedded more densely than anomalies, with a continuous trend toward the anomalous region.
The Jacobian analysis suggests that the latent states are linearly independent, as indicated by the full-rank property observed in the Singular Value Decomposition (SVD) analysis. However, this does not inherently enforce an orthonormal projection, meaning that scale freedom exists and angular relationships are not necessarily preserved.
Nevertheless, from a phase-space perspective, the qualitative relationship between states remains intact. Anomalous states continue to map to regions outside the bounded domain of normal states, ensuring that the essential structural distinction is preserved, even if exact angles and scales are distorted.
Our objective is to develop methods that integrate physics-inspired consistency principles while imposing only the necessary constraints. This approach maintains the model’s flexibility and ability to learn effectively. Therefore, we do not enforce an explicit orthonormality condition in the AE design.
Figure 3: Latent node state–derivative pairs for nodes (2, 6) are shown in panels (a) and (b), and for nodes (3, 7) in panels (d) and (e). The corresponding phase-space representations for these pairs are displayed in panels (c) and (f), respectively. In all phase-space plots, the node values have been smoothed using a sliding average window of 3.
Let’s have a more detailed view on the phase space evolution. Fig. 2a illustrates a clear drift toward the anomalous region, highlighting a gradual deviation in system behavior. Fig. 2b presents the evolution of $\rho$, which quantifies the causal approximation of state-derivative pairs. In the normal range, $\rho$ remains stable on average, indicating a consistent dynamical description. However, as the system enters the anomalous regime, a significant increase in $\rho$ is observed, suggesting a loss in the dynamical consistency of the system. The theoretical derivation showed that when system dynamics deviate from a stable state, the increasing complexity requires a higher-dimensional embedding to adequately capture the system’s behavior. We hypothesized that this effect would lead to the collapse of state-derivative approximations, serving as an indicator of anomalous behavior. Our observations now confirm this, as we see a loss of dynamical consistency in the latent space, reinforcing our hypothesis that anomalies emerge when the system transitions beyond the representational capacity of the learned embedding.
In Fig. 2c, we observe the evolution of the phase space representation over time, transitioning from normal to abnormal behavior. Initially, the phase space exhibits an attractor-like structure around which the dynamics evolve in a stable manner. As the system progresses, we observe increased dispersion and a gradual drift away from the attractor toward the anomalous region. While in the range characterized as normal, the state-derivative approximation remains stable, as seen in Fig. 2a. However, subtle deviations from the attractor’s dynamics can already be detected in the earliest phase of the transition. This aligns with our second hypothesized effect, which suggests that even minor disruptions in the system’s stability could serve as an early indicator of anomalies.
These findings reinforce that our model is not merely a data-driven classifier for anomaly detection but instead captures meaningful insights into the underlying system dynamics.
IV-C Comparison with Literature Benchmarks
TABLE II: Performance comparison of different approaches benchmarks for anomaly detection in the C-MAPSS dataset.
Algorithm Acc. (%) Prec. (%) Rec. (%) F1 (%)
TDC-AE $\mathbf{99.3}\pm 0.3$ $\mathbf{99.99}\pm 0.03$ $\mathbf{98.30}\pm 0.07$ $\mathbf{99.14}\pm 0.04$
DeepLSTM-AE [19] 96.45 94.81 98.12 96.44
Random Forest [19] 92.33 97.40 86.66 91.71
CNN-AE [19] 92.22 89.08 95.86 92.35
SMOTE Transformer [7] 89.59 93.95 91.77 -
XGBoost [19] 84.57 77.73 96.00 85.90
Dense-AE [18] - 89.6 72.4 80.1
We have discussed the consistency of our method and analyzed how the TDC-AE aligns with the mathematical foundation developed in this paper. Now, we evaluate its detection performance in comparison to literature benchmarks.
Tab. II presents a comparison across different architectures. Our approach outperforms all reported models across all detection metrics, demonstrating consistent results across repeated attempts. The latent node threshold in all attempts was set using the 75th percentile of the validation set.
Even complex models, such as LSTM-AEs, Transformer with attention mechanism designed to capture long-range temporal correlations, do not achieve superior performance. In [19], the exact LSTM architecture is not explicitly provided, so we base our assumptions on similar studies of complex physical systems [9], considering the following configuration: sequence length $L=48$, hidden dimension $d_{\text{hidden}}=16$, and input dimension $d_{\text{input}}=24$. This results in a Multiply-Accumulate Operations (MAC) count of 245,760.
In contrast, our approach achieves superior detection performance with significantly lower computational complexity, requiring only 2,688 MACs.
Furthermore, a variant of the TDC-based architecture has also been applied to the BATADAL dataset, which models cyber-physical water distribution systems. This approach outperformed the challenge-winning method [31], which is grounded in explicit physical modeling. These results underscore the potential of the proposed framework to generalize across structurally different dynamical systems and application domains [32].
V Conclusion
We introduced an unsupervised framework for anomaly detection in complex physical systems, grounded in classical embedding theory and physics-inspired consistency principles, particularly state-derivative relations in the embedding space. To translate this theoretical foundation into practice, we developed TDC-AE, an algorithm designed to capture system dynamics efficiently.
The results demonstrate that our chosen embedding strategy can effectively capture time correlation from a single time step. TDC-AE outperforms benchmarks on the C-MAPSS dataset, surpassing computationally intensive models like LSTMs and Transformers, while achieving a nearly 200x reduction in MAC operations, keeping them below 3,000 MACs. This efficiency makes our approach particularly suitable for lightweight edge computing.
Appendix A Theoretical Background
A-A General Definition of an Oscillatory Dynamical System
Figure 4: A driven damped pendulum with angular displacement $\theta$, velocity $\dot{\theta}$, damping $-\gamma\dot{\theta}$, and external driving force $F$.
A general oscillatory dynamical system is governed by a set of differential equations describing periodic motion. In its most general form, the system evolves according to
$\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x},t),$ (5)
where $\mathbf{x}\in\mathbb{R}^{n}$ represents the state variables, and $\mathbf{f}(\mathbf{x},t)$ defines the governing dynamics. A common class of oscillatory systems follows a second-order differential equation of the form
$\ddot{\theta}+g(\theta,\dot{\theta})+h(\theta,t)=0,$ (6)
where $g(\theta,\dot{\theta})$ represents dissipative forces, such as damping, and $h(\theta,t)$ accounts for external periodic forcing. A widely studied case is the damped driven oscillator, as shown in Fig. 4, which satisfies the equation
$\ddot{\theta}+\gamma\dot{\theta}+\omega_{0}^{2}\sin\theta=A\cos(\omega_{\text{drive}}t),$ (7)
where $\theta$ denotes the angular displacement and $\dot{\theta}$ its corresponding angular velocity. The parameters $\gamma$ and $\omega_{0}$ represent the damping coefficient and the natural frequency of the system, respectively, while $A$ and $\omega_{\text{drive}}$ define the amplitude and frequency of an external periodic driving force [33]. To analyze deviations from the nominal oscillatory state, we introduce a slow drift in both the angular displacement and velocity components by modifying the evolution equations as
${\theta}\leftarrow{\theta}+\alpha t,\quad\dot{\theta}\leftarrow\dot{\theta}+\beta t.$ (8)
For illustration, we consider a system with parameters $\gamma=0.2$, $\omega_{0}=1.0$, $A=0.8$, $\omega_{\text{drive}}=1.2$, and drift rates $\alpha=0.005$ and $\beta=0.002$. Under these conditions, the system initially remains within a bounded oscillatory regime but gradually exhibits deviations due to the drift terms, leading to an eventual departure from the confined state space. The effect of this perturbation is visualized in Figure 1, where the phase space trajectory initially remains within a bounded region but progressively transitions into an unbounded state.
Appendix B Implementation details and numerical methods
B-A The Box-Counting Dimension in a dammped driven Pendulum
To illustrate the concept of fractal dimension, we numerically estimated the box-counting dimension for a slice of the phase space of the damped driven pendulum. We considered the first 100–1000 time steps and applied the box-counting method to a subset of the $(\theta,\dot{\omega})$ plane.
Using a range of box sizes $\epsilon$, we counted the number of occupied boxes $N(\epsilon)$ that contained at least one trajectory point. The chosen values of $\epsilon$ were logarithmically spaced as:
$\epsilon\in\{10^{-2},10^{-1.9},10^{-1.8},\dots,10^{-0.4}\}.$
These values were selected to balance resolution constraints and ensure a meaningful estimation within the given simulation timestep.
The relationship between $\log N(\epsilon)$ and $\log(1/\epsilon)$ was then analyzed through a linear fit, where the slope provides an estimate of the box-counting dimension.
Fig. 5a illustrates the fitting result, and Fig. 5b shows the selected phase-space slice used for the computation.
Figure 5: (a) Log-log plot of the box-counting method, showing the linear fit used to estimate the fractal dimension. (b) Phase-space slice of the damped driven pendulum used for the computation. The schematic shows 48 grid cells with $\epsilon\approx 0.45$. The roughly corresponding value pair is highlighted as a green square in panel (a).
B-B TDC-AE
Algorithm 1 Temporal Differential Consistency Loss Enhanced Autoencoder (TDC-AE)
1:  Initialize Autoencoder with input $X_{t}\in\mathbb{R}^{k}$ and latent variables and time derivatives ($\mathbf{z},\dot{\mathbf{z}})\in\mathbb{R}^{n}$, Optimizer, MSE Loss
2:  for each epoch in training_epochs do
3:     for each batch in training_data do
4:        Perform forward pass through the Autoencoder and Encoder: $X_{\text{rec}}\leftarrow\text{Autoencoder}(X_{t})$
5:        Compute latent representations for neighboring time steps: $(\mathbf{z}_{t-1},\dot{\mathbf{z}}_{t-1})\leftarrow\text{Encoder}(X_{t-1})$ $(\mathbf{z}_{t+1},\dot{\mathbf{z}}_{t+1})\leftarrow\text{Encoder}(X_{t+1})$
6:        Compute central difference derivative: $\Delta_{t}\mathbf{z}\leftarrow(\mathbf{z}_{t+1}-\mathbf{z}_{t-1})/2\Delta_{t}$
[…]
10:        Update the Autoencoder parameters using the optimizer
11:     end for
12:  end for
B-C Implementation of the TDC-AE and Hardware
TABLE III: Hyperparameters used in training the autoencoder for the C-MAPSS (FD001) dataset.
Hyperparameter Value
Optimizer Adamax
Activation Function Tanh
Number of Dense Layers 6
Architecture 24-24-24-8-24-24-24
Learning Rate 0.003
Batch Size 32
$\alpha$ 100
The experiments were conducted on a system running Windows 11, powered by a 13th Gen Intel Core i7-13700H processor (2.4 GHz, 14 cores, 20 threads), along with 32 GB of RAM and an NVIDIA GeForce RTX 4070 Laptop GPU.
The model was implemented using PyTorch 2.3.0 with CUDA 11.8, leveraging GPU acceleration to enhance computational efficiency and reduce training time. The autoencoder’s architecture and training setup are detailed in Table III.
Appendix C Figures
Figure 6: The two loss terms used in TDC-informed training across five independent attempts on the C-MAPSS test dataset.
Figure 7: Schematic of temporal difference consistency in the latent space of an autoencoder. The derivative node $z_{\dot{}}$ is aligned with the central difference of the static node $z$ using time steps $t-1$, $t$, and $t+1$.